OpenStudy (anonymous):

Given quadrilateral RSTU, determine if each pair of sides (if any) are parallel and which are perpendicular for the coordinates of the vertices. R(0, 0), S(6, 3), T(5, 5), U(-1, 2)

OpenStudy (anonymous):

Medals 0 * slope 2 = slope1 => parallel * slope 2=-1/slope1 => perpendicular slope(RS)= (y2-y2)/(x2-x1)=(3-0)/6-0)=1/2 slope(RT)=(5-0)/(5-0)=1 slope(RU)=(2-0)/(-1-0)=2 slope(ST)=(5-3)(5-6)=2/-1=-2 slope(SU)=(2-3)(-1-6)=1/-7=-1/7 slope(TU)=(5-2)(5-(-1))=3/6=1/2

OpenStudy (anonymous):

slope(RS)=1/2=slope(TU) => they are parallel slope(ST)=-2 it is perpendicular to both RS and TU slope 2=-1/slope(RS)=-1/(1/2)=-2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

YW